The length of a rectangle is 3 feet more than the width. The area of the rectangle is 54 square feet. Find the dimensions of the rectangle.
Answer by checkley77(12844)

L=W+3
LW=54
(W+3)W=54
W^2+3W-54=0
(W+9)(W-6)=0
W-6=0
W=6 ANS.
L=6+3=9 ANS.
PROOF:
6*9=54
54=54